OpenStudy (anonymous):

I need help with the problem, 4 sqrt(u) (sqrt(6) + 7sqrt(24)) Every time I do it I'm somehow wrong...

OpenStudy (anonymous):

is that really a \(u\) in the first term?

OpenStudy (anonymous):

\[4\sqrt{u}(\sqrt{6} + 7\sqrt{24})\]

OpenStudy (anonymous):

oh , so it is a \(u\)

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

since \(24=4\times 6\) you have \(\sqrt{24}=\sqrt{4\times 6}=\sqrt{4}\sqrt{6}=2\sqrt{6}\) making this \[4\sqrt{u}(\sqrt{6} + 14\sqrt{6})\]

OpenStudy (anonymous):

then combine like terms, which gives you \[4\sqrt{u}\times 15\sqrt{6}\]

OpenStudy (anonymous):

I've been able to get that far but after that I somehow mess up...

OpenStudy (anonymous):

final answer would be \(60\sqrt{6u}\)

OpenStudy (anonymous):

Ohh I forgot about the 1 before the sqrt(6)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

the one? yw